# Screened Thin-Target Bremsstrahlung with Partially-Ionized High-Z Species

### v1

@@ III.2 Screened Cross Section
$I_{2}(\bar{b})$ $=-\frac{\alpha Z^{2}r_{0}^{2}}{2\pi\bar{k}\,\bar{\bm{p}_{0}}}\Biggl\{\frac{2\bar{p}\left(4\bar{E}_{0}^{2}+\bar{b}^{2}\right)}{W^{2}}\left(\bar{E}_{0}\frac{D_{1}}{\bar{k}}-2+\bar{b}^{2}\right)$
$+\frac{L_{1}}{W}\Bigl[\bar{k}(D_{1}+2\bar{b}^{2})$ (80)
$\quad+\frac{2\bar{k}}{D_{1}}\bigl(\bar{b}^{4}+2\bar{b}^{2}(\bar{E}_{0}^{2}+\bar{p}^{2})-8\bar{E}_{0}\bar{E}\bigr)\Bigr]$
$+\frac{4\bar{E}_{1}^{2}+\bar{b}^{2}}{\bar{k}W^{2}}\Bigl(D_{1}^{2}+2(\bar{E}_{0}\bar{E}-1)D_{1}$ (81)
$\quad+\bar{b}^{2}(D_{1}-2\bar{E}\bar{k})\Bigr)$
$+\frac{\bar{k}^{2}L_{2}}{D_{1}\,\|\bar{\bm{p}_{0}}-\bar{\bm{k}}\|}\Biggl[\frac{2}{D_{1}}\bigl(4\bar{E}^{2}+\bar{b}^{2}(1-D_{1})\bigr)$
$\quad+\frac{D_{1}^{2}+2(\bar{E}_{0}\bar{E}-1)D_{1}+\bar{b}^{2}(D_{1}-2\bar{E}\bar{k})}{2\|\bar{\bm{p}_{0}}-\bar{\bm{k}}\|^{2}}\Biggr]$
$-\frac{4\bar{k}}{D_{1}}\,\bar{b}^{2}\ln(\bar{E}+\bar{p})\Biggr\}.$ (82)
@@
$D_{1}$ $=2\bar{k}\,\bigl(\bar{E}_{0}-\bar{\bm{p}_{0}}\cos\theta_{1}\bigr)$ (83)
$W$ $=\Biggl[\left(\bar{\bm{p}_{0}}\,\frac{D_{1}}{\bar{k}}\right)^{2}+2\bar{b}^{2}\left(\bar{E}_{0}\frac{D_{1}}{\bar{k}}-2\right)+\bar{b}^{4}\Biggr]^{1/2}$ (84)
$L_{1}$ $=\ln\frac{(\bar{E}_{0}\bar{E}-1)\,D_{1}/\bar{k}+\bar{E}\,\bar{b}^{2}+\bar{p}\,W}{(\bar{E}_{0}\bar{E}-1)\,D_{1}/\bar{k}+\bar{E}\,\bar{b}^{2}-\bar{p}\,W}$ (85)
$L_{2}$ $=\ln\frac{(\bigl\|\bar{\bm{p}_{0}}-\bar{\bm{k}}\bigr\|+\bar{p})^{2}+\bar{b}^{2}}{(\bigl\|\bar{\bm{p}_{0}}-\bar{\bm{k}}\bigr\|-\bar{p})^{2}+\bar{b}^{2}}.$ (86)
@@ IV Cross Section Results
The equations (15), (19), and (57) have been implemented in a Matlab program [16]. While Eqs. (15) and (57) can be evaluated directly from their analytical forms,
Figure 1: Comparison of the DDCS obtained from the closed form Eq. (57) (solid lines) and from a numerical integration of Eq. (39) (markers), for several ionized states of copper, and two Yukawa exponentials.
the numerical integration of the RDP cross section requires careful handling of hypergeometric functions and is computationally more demanding.
To assess the consistency of the analytical DDCS for partially ionized species, results from the closed-form expression (57) are compared with those obtained by direct numerical integration of Eq. (39) for all ionization states of a given atom. Fig. 1 shows an excellent agreement for several ionization degrees of aluminum ($Z=13$), where two Yukawa exponentials have been used for the description of the potential. As expected, the emission increases with the ionization state, i.e. with the effective ion charge $Z_{s^{\prime}}$. The fully stripped limit is Sauter’s expression Eq. (15) for a pure Coulomb field. The neutral case $Z_{0}$ corresponds by definition to the expression found by Haug [17], except that in the present model the number of exponentials is $Z$-dependent and not fixed to three, allowing a better fitting of the atomic potential.
@@ IV.1 Multi-Yukawa and Ionized Cross Sections
We begin by examining how the atomic form factor (AFF) is influenced by the number of exponentials used in the multi-Yukawa model. The top panel of Fig. 2 shows the AFF as a function of the nuclear recoil momentum for a neutral copper atom, calculated with several values of $N_{Y}$ and compared with the “exact” form factor obtained from density-functional theory (DFT) [54, 50] using the Gaussian code [12]. The bottom panel of Fig. 2 illustrates the corresponding impact on the doubly differential cross section (DDCS). The incident-electron energy is set to 1.7 MeV, and two small emission angles, $0^{\circ}$ and $5^{\circ}$, are considered, since screening effects are most pronounced in the forward direction. As expected, the DDCS increases slightly with $N_{Y}$, because the screening factor for $N_{Y}=1$ is larger than for $N_{Y}=2,3$ at low values of $\bar{q}$.
The DDCS is calculated for various ionized states of gold ($Z=79$) and shown in Fig. 3. Although the cross section is monotonically increasing with the ion charge $Z_{s^{\prime}}$ for low values of $k$, a minimum is found at non-neutral states for higher photon energies.
The non-monotonic behavior of the doubly differential cross section with increasing ionization state at higher photon energies $k$ originates from the atomic form factor, which reflects the effective nuclear charge experienced during the interaction. At higher photon energies, the form factor does not vary monotonically with $Z_{s^{\prime}}$, as illustrated in Fig. 4, where $F_{Z_{s^{\prime}}}$ is plotted as a function of the recoil momentum $\bar{q}(k)$ for several ionization states.
@@
Since the atomic form factor is the Fourier transform of the electronic density $n_{s^{\prime}}(\mathbf{r})$,
$F_{Z_{s^{\prime}}}(\bar{q})$ $=\frac{1}{Z}\int_{\mathbb{R}^{3}}n_{s^{\prime}}(\mathbf{r})\,e^{i\bar{\bm{q}}\cdot\mathbf{r}}\,d^{3}r$ (87)
$=1-\sum_{i}\bar{A}_{i}\,\frac{\bar{q}^{2}+q_{Z_{s^{\prime}}}\bar{b}_{i}^{2}}{\bar{q}^{2}+\bar{b}_{i}^{2}},$ (88)
the relative behavior of the form factors for different ionization states is strongly influenced by the corresponding radial electron densities around the nucleus. When these densities cross, the form factors naturally exhibit similar non-monotonic behavior.
@@
The intersection radius between two radial bound-electron density profiles, $n_{s^{\prime}}(r)$ and $n_{s^{\prime\prime}}(r)$, provides a convenient way to visualize how ionization modifies the spatial extent of the electronic cloud. If the bound-electron distribution contracted strictly monotonically with increasing ionization, no radial crossings would occur. Each more-ionized state would be confined to a smaller radius. Conversely, a radial crossing between $n_{s^{\prime}}$ and $n_{s^{\prime\prime}}$ indicates that upon removing electrons, the residual bound charge has redistributed onto a larger characteristic shell at a certain radius, so that the screening length is increased rather than decreased. Among the possible intersections, the outermost crossing is the most physically relevant, as it occurs in the outer region of the atom where the residual bound-electron density dominates and thus determines the long-range screening behavior (low $\bar{q}$-domain). The presence of such crossings therefore reveals that the contraction of the electronic cloud with ionization is not perfectly monotonic: the residual electrons undergo redistribution which in turn lead to the non-monotonic behavior of the atomic form factor $F_{Z_{s^{\prime}}}(\bar{q})$ with ionization, as observed in Fig. 4.
@@ IV.2 Comparison with Data for the Doubly Differential Cross Section
In Figs. 6, 7, 8, the solid red lines represent the Sauter formula, Eq. (15); the dashed green curves correspond to the multi-Yukawa Born approximation, Eq. (57); and the dashed blue curves show the integrated RDP result, Eq. (19). The total cross section obtained from Eq. (3) is plotted as a dashed black line. The RDP cross section was computed by numerically integrating the TDCS Eq. (19), over the scattering angles $(\theta,\phi)$ using an adaptive grid method that ensures a relative accuracy better than $10^{-6}$ [16]. Measurement uncertainties include statistical errors (error bars), which affect both the spectral shape and the absolute magnitude of the cross sections, and systematic errors, which influence only the absolute magnitude [47, 42]. The contours represent the combined statistical and systematic uncertainties.
@@
$\frac{\alpha Z_{s}}{\beta_{0}},\;\frac{\alpha Z_{s}}{\beta}\ll 1,$ (89)
@@
is no longer satisfied, and of the FSM wave-function validity. [^2] Note that Furry–Sommerfeld–Maue condition (31) involves the local angle $\theta_{pr}$ between the electron momentum and its position vector in the Coulomb field, rather than the measured photon emission angle $\theta_{0}$. Although these two quantities are distinct, they are correlated through the recoil momentum $\bar{\bm{q}}=\bar{\bm{p}}_{0}-\bar{\bm{p}}-\bar{\bm{k}}$ as larger $\bar{q}$ corresponds to smaller effective electron–nucleus separations and therefore to stronger local deflections [18]. The FSM approximation is expected to deteriorate in regions of large $\bar{q}$, which typically occur for large emission angles. The breakdown of the FSM condition is likely the dominant cause, given that Eq. (89) contains no angular dependence, and the lower-left panel ($Z=79$, $\theta_{0}=10^{\circ}$) does not reveal any obvious failure of the Born-based cross sections (Sauter and screened Bethe–Heitler).
@@
At larger angles, where screening effects are weaker, the Sauter and Bethe–Heitler results are expected to converge towards each other, since the screening correction within the Born approximation becomes small. In this regime, the dominant contribution should therefore arise from the RDP cross section, which lies significantly below the experimental values for $\theta_{0}=60^{\circ}$, consistent with the breakdown of the FSM approximation in the high-$\bar{q}$ region. It is however important to note that the corresponding cross sections are two orders of magnitude lower than those obtained for emissions at $10^{\circ}$.
@@
The onset of this behavior also appears at smaller emission angles for higher incident electron energies, as shown by comparison with data in Figs. 7 and 8. The validity of the FSM condition (31) becomes increasingly restricted at higher incident energies, since the momentum transferred to the nucleus $\bar{q}$ rises for a given emission angle, probing smaller distances from the nucleus according to $r\sim 1/\bar{q}$ [18]. As the relevant interaction radius decreases, the local deflection $\vartheta_{pr}$ of the electron in the Coulomb field grows, and the FSM approximation can fail even for comparatively small observable angles. This trend is evident in Fig. 7 for 4.54 MeV electrons incident on gold ($Z=79$), where a noticeable discrepancy between the model and the experimental measurements already appears at $\theta_{0}\simeq 6^{\circ}$. At such angles, however, the corresponding emission intensity is an order of magnitude lower than in the purely forward direction ($\theta_{0}=0^{\circ}$). As the emission angle increases, all theoretical components (RDP, Sauter, and BHs) shift downward relative to the experimental data, such that their combination, Eq. (3), lies systematically below the measurements, even though these angles remain smaller than in Fig. 6.
@@
At these higher energies, an additional source of uncertainty arises from the photomultiplier response function used by Starfelt and Koch, which was obtained through Monte Carlo simulations performed by other authors [47]. Nevertheless, the systematic downward shift of all theoretical components strongly indicates that the principal cause of the discrepancy is the breakdown of the FSM approximation in the high-$\bar{q}$ regime, where the local perturbative expansion ceases to hold.
Finally, as noted by Roche, Ducos, and Proriol [43], the RDP model remains accurate up to electron energies of approximately 50 MeV. The present work extends their formulation by incorporating screening effects into the cross-section calculation. Figure 10 (see Appendix B) presents bremsstrahlung cross sections as a function of photon energy for an incident electron energy of 30 MeV. Unfortunately, no experimental data exist above about 10 MeV with which these predictions can be compared. The bremsstrahlung community would therefore greatly benefit from new measurements in this higher-energy range, both to expand the accessible parameter space and to better assess the accuracy of current theoretical models.

### v2

@@ III.2 Screened Cross Section
$I_{2}({b})$ $=-\frac{\alpha Z^{2}r_{0}^{2}}{2\pi{k}\,{{p}_{0}}}\Biggl\{\frac{2{p}\left(4{E}_{0}^{2}+{b}^{2}\right)}{W^{2}}\left({E}_{0}\frac{D_{1}}{{k}}-2+{b}^{2}\right)$
$+\frac{L_{1}}{W}\Bigl[{k}(D_{1}+2{b}^{2})$ (72)
$\quad+\frac{2{k}}{D_{1}}\bigl({b}^{4}+2{b}^{2}({E}_{0}^{2}+{p}^{2})-8{E}_{0}{E}\bigr)\Bigr]$
$+\frac{4{E}_{1}^{2}+{b}^{2}}{{k}W^{2}}\Bigl(D_{1}^{2}+2({E}_{0}{E}-1)D_{1}$ (73)
$\quad+{b}^{2}(D_{1}-2{E}{k})\Bigr)$
$+\frac{{k}^{2}L_{2}}{D_{1}\,\|{\bm{p}_{0}}-{\bm{k}}\|}\Biggl[\frac{2}{D_{1}}\bigl(4{E}^{2}+{b}^{2}(1-D_{1})\bigr)$
$\quad+\frac{D_{1}^{2}+2({E}_{0}{E}-1)D_{1}+{b}^{2}(D_{1}-2{E}{k})}{2\|{\bm{p}_{0}}-{\bm{k}}\|^{2}}\Biggr]$
$-\frac{4{k}}{D_{1}}\,{b}^{2}\ln({E}+{p})\Biggr\}.$ (74)
@@
$D_{1}$ $=2{k}\,\bigl({E}_{0}-{{p}_{0}}\cos\theta_{1}\bigr)$ (75)
$W$ $=\Biggl[\left({{p}_{0}}\,\frac{D_{1}}{{k}}\right)^{2}+2{b}^{2}\left({E}_{0}\frac{D_{1}}{{k}}-2\right)+{b}^{4}\Biggr]^{1/2}$ (76)
$L_{1}$ $=\ln\frac{({E}_{0}{E}-1)\,D_{1}/{k}+{E}\,{b}^{2}+{p}\,W}{({E}_{0}{E}-1)\,D_{1}/{k}+{E}\,{b}^{2}-{p}\,W}$ (77)
$L_{2}$ $=\ln\frac{(\bigl\|{\bm{p}_{0}}-{\bm{k}}\bigr\|+{p})^{2}+{b}^{2}}{(\bigl\|{\bm{p}_{0}}-{\bm{k}}\bigr\|-{p})^{2}+{b}^{2}}.$ (78)
@@ IV Cross Section Results
The equations (12), (16), and (49) have been implemented in a Matlab program [16]. While Eqs. (12) and (49) can be evaluated directly from their analytical forms, the RDP contribution is evaluated by numerically integrating the TDCS in Eq. (16) over the scattering angles $(\theta,\phi)$ using an adaptive grid method, achieving a relative accuracy better than $10^{-6}$ [16].
Figure 1: Left: Region of the parameter space where the screening is prominent, namely soft-photons and forward emissions, for an incoming electron of energy $E_{0}=4.54$ MeV on a gold target. Right: Region of the parameter space where Coulomb corrections are required, i.e. hard photons and forward emissions. In the ratios, $d^{2}\sigma$ stands for $d^{2}\sigma/dkd\Omega_{k}$.
Figure 1 provides physical motivation for the separation of screening and Coulomb contributions embodied in the Olsen–Maximon–Wergeland (OMW) formulation, Eq. (3). The left panel shows the logarithm of the ratio between the unscreened (Sauter) reference and the corresponding screened Bethe–Heitler (multi-Yukawa) DDCS, as a function of the emitted photon energy and emission angle, for an incident electron energy of $E_{0}=4.54$ MeV. This representation highlights the importance of screening effects, which are most pronounced in the soft-photon, forward-emission domain. The right panel displays the logarithm of the ratio of the leading-order RDP (Elwert–Haug) cross section to Sauter’s Coulomb-field result, illustrating the onset of Coulomb (nuclear) corrections toward forward angles and higher photon energies. Because the next-to-leading-order RDP expression is not reliable in the high-frequency (tip) region, the leading-order result is used in the ratio shown. [^2] The next-to-leading-order RDP theory was derived under the assumption that both the incident and scattered electron energies are large, and therefore loses accuracy in the high-frequency (tip) region.
Figure 1 is intended to illustrate the physical motivation for treating the screening correction additively in Eq. (3); a detailed assessment of the OMW additivity rule can be found in Ref. [20].
Figure 2: Comparison of the DDCS obtained from the closed form Eq. (49) (solid lines) and from a numerical integration of Eq. (32) (markers), for several ionized states of aluminum, and two Yukawa exponentials.
To assess the consistency of the analytical DDCS for partially ionized species, results from the closed-form expression (49) are compared with those obtained by direct numerical integration of Eq. (32) for all ionization states of a given atom. Fig. 2 shows an excellent agreement for several ionization degrees of aluminum ($Z=13$), where two Yukawa exponentials have been used for the description of the potential. As expected, in the screening-dominant (large wavelength) region, the emission increases with the ionization state, i.e. with the effective ion charge $Z_{s}$. The fully stripped limit is Sauter’s expression Eq. (12) for a pure Coulomb field. The neutral case $Z_{0}$ corresponds by definition to the expression found by Haug [18], except that in the present model the number of exponentials is $Z$-dependent and not fixed to three, allowing a better fitting of the atomic potential.
@@ IV.1 Multi-Yukawa and Ionized Cross Sections
We begin by examining how the atomic form factor (AFF) is influenced by the number of exponentials, $N_{Y}$, used in the multi-Yukawa model. The top panel of Fig. 3 shows the AFF as a function of the nuclear recoil momentum for a neutral copper atom, calculated with several values of $N_{Y}$ and compared with the “exact” form factor obtained from density-functional theory (DFT) [55, 48] using the Gaussian code [12]. The bottom panel of Fig. 3 illustrates the corresponding impact on the doubly differential cross section (DDCS). The incident-electron energy is set to 1.7 MeV, and two small emission angles, $0^{\circ}$ and $5^{\circ}$, are considered, since screening effects are most pronounced in the forward direction. As expected, the DDCS increases slightly with $N_{Y}$, because the screening factor for $N_{Y}=1$ is larger than for $N_{Y}=2,3$ at low values of $q$.
The DDCS is calculated for various ionized states of gold ($Z=79$) and is shown in Fig. 4. Although the cross section is monotonically increasing with the ion charge $Z_{s}$ for low values of $k$, a minimum is found at non-neutral states for higher photon energies.
@@
The non-monotonic behavior of the doubly differential cross section with increasing ionization state at higher photon energies $k$ originates from the atomic form factor, which reflects the effective nuclear charge experienced during the interaction. At higher photon energies, the form factor does not vary monotonically with $Z_{s}$, as illustrated in Fig. 5, where $F_{Z_{s}}$ is plotted as a function of the recoil momentum $q(k)$ for several ionization states.
Since the atomic form factor is the Fourier transform of the spatial electronic density $n_{s}$,
$F_{Z_{s}}({q})$ $=\frac{1}{Z}\int_{\mathbb{R}^{3}}n_{s}(\mathbf{r})\,e^{i{\bm{q}}\cdot\mathbf{r}}\,d^{3}r$ (79)
$=1-\sum_{i}{A}_{i}\,\frac{{q}^{2}+q_{Z_{s}}{b}_{i}^{2}}{{q}^{2}+{b}_{i}^{2}},$ (80)
where $\bm{r}$ is in units of the reduced Compton wavelength and $n_{s}$ is rescaled accordingly so as to preserve its normalization, the relative behavior of the form factors for different ionization states is strongly influenced by the corresponding radial electron densities around the nucleus. When these densities cross, the form factors can exhibit similar non-monotonic behavior.
@@
The intersection radius between two radial bound-electron density profiles, $n_{s}(r)$ and $n_{s^{\prime}}(r)$, provides a convenient way to visualize how ionization modifies the spatial extent of the electronic cloud. If the bound-electron distribution contracted strictly monotonically with increasing ionization, no radial crossings would occur. Each more-ionized state would be confined to a smaller radius. Conversely, a radial crossing between $n_{s}$ and $n_{s^{\prime}}$ indicates that upon removing electrons, the residual bound charge has redistributed onto a larger characteristic shell at a certain radius, so that the screening length is increased rather than decreased. Among the possible intersections, the outermost crossing is the most physically relevant, as it occurs in the outer region of the atom where the residual bound-electron density dominates and thus determines the long-range screening behavior (low $q$-domain). The presence of such crossings therefore indicates that the contraction of the electronic cloud with ionization is not perfectly monotonic: the residual electrons undergo redistribution which in turn leads to the non-monotonic behavior of the atomic form factor $F_{Z_{s}}(q)$ with ionization, as observed in Fig. 5.
@@ IV.2 Comparison with Data for the Doubly Differential Cross Section
In Figs. 7, 8, 9, the solid red lines represent the Sauter formula, Eq. (12); the dashed green curves correspond to the multi-Yukawa Born approximation, Eq. (49); and the dashed blue curves show the integrated RDP result, Eq. (16). The total cross section obtained from Eq. (3) is plotted as a dashed black line. Measurement uncertainties include statistical errors (error bars), which affect both the spectral shape and the absolute magnitude of the cross sections, and systematic errors, which influence only the absolute magnitude [51, 43]. The contours represent the combined statistical and systematic uncertainties.
@@
$\frac{\alpha Z_{s}}{\beta_{0}},\;\frac{\alpha Z_{s}}{\beta}\ll 1$ (81)
@@
is no longer satisfied, and of the FSM wave-function validity. The breakdown of the FSM condition is likely the dominant cause, given that Eq. (81) contains no angular dependence, and the lower-left panel ($Z=79$, $\theta_{0}=10^{\circ}$) does not reveal any obvious failure of the Born-based cross sections (Sauter and screened Bethe–Heitler).
@@
At larger angles, where screening effects are weaker, the Sauter and Bethe–Heitler results are expected to approach each other, since the screening correction within the Born approximation becomes small. In this regime, the dominant contribution should therefore arise from the RDP cross section, which lies significantly below the experimental values for $\theta_{0}=60^{\circ}$, consistent with the breakdown of the FSM approximation in the high-$q$ region. [^4] Note that Furry–Sommerfeld–Maue condition (26) involves the local angle $\vartheta_{pr}$ between the electron momentum and its position vector in the Coulomb field, rather than the measured photon emission angle $\theta_{0}$. Although these two quantities are distinct, they are correlated through the recoil momentum ${\bm{q}}={\bm{p}}_{0}-{\bm{p}}-{\bm{k}}$ as larger q{q} corresponds to smaller effective electron–nucleus separations and therefore to stronger local deflections [17]. The FSM approximation is expected to deteriorate in regions of large q{q}, which typically occur for large emission angles. It is however important to note that the corresponding cross sections are two orders of magnitude lower than those obtained for emissions at $10^{\circ}$.
@@
The onset of this behavior also appears at smaller emission angles for higher incident electron energies, as shown by comparison with data in Figs. 8 and 9. The validity of the FSM condition (26) becomes increasingly restricted at higher incident energies, since the momentum transferred to the nucleus, q{q}, rises for a given emission angle, probing smaller distances from the nucleus according to $r\sim 1/{q}$ [17]. As the relevant interaction radius decreases, the local deflection $\vartheta_{pr}$ of the electron in the Coulomb field grows, and the FSM approximation can fail even for comparatively small observable angles. This trend is evident in Fig. 8 for 4.54 MeV electrons incident on gold ($Z=79$), where a noticeable discrepancy between the model and the experimental measurements already appears at $\theta_{0}\simeq 6^{\circ}$. At such angles, however, the corresponding emission intensity is an order of magnitude lower than in the purely forward direction ($\theta_{0}=0^{\circ}$). It is important to note that in the near-forward direction $(\theta_{0}\sim 0^{\circ})$, the RDP cross section decreases sharply in the high-frequency limit, in a manner similar to the Born results (see the two upper panels of Fig. 8). Although this behavior is mitigated as the emission angle increases, the RDP cross section cannot be expected to be accurate in the tip region, since the next-to-leading-order correction relies on the assumption that the scattered electron remains highly relativistic. By contrast, the leading-order (EH) cross section does not exhibit such a decrease in the tip region, although it is not shown here. As the emission angle increases, all theoretical components (RDP, Sauter, and BHs) shift downward relative to the experimental data, such that their combination, Eq. (3), lies systematically below the measurements, even though these angles remain smaller than in Fig. 7.
@@
At these higher energies, an additional source of uncertainty arises from the photomultiplier response function used by Starfelt and Koch, which was obtained through Monte Carlo simulations performed by other authors [51]. Nevertheless, the systematic downward shift of all theoretical components strongly indicates that the principal cause of the discrepancy is the breakdown of the FSM approximation in the high-$q$ regime, where the local perturbative expansion ceases to hold.
Roche, Ducos, and Proriol originally indicated that their formulation remains applicable up to electron energies of approximately 50 MeV [45]. In the present work, this framework is employed in its consistently ordered form (see Eq. (20)). Figure 10 shows the components of the model in Eq. (3) for an incident electron energy of 30 MeV on a gold target. The various contributions behave consistently and exhibit no qualitative discrepancies relative to the trends observed at lower energies. The choice of 30 MeV corresponds to approximately three times the highest incident energy for which experimental data are currently available. These results suggest that the present formulation can reasonably be applied at such energies, provided the system remains in a regime where the breakdown of the FSM approximation is limited. When the FSM approximation is no longer adequate, an exact treatment of the Coulomb field using Dirac wave functions provides a natural alternative, assuming that screening corrections can still be incorporated additively within the OMW framework.
Figure 10: Top: Theoretical model components for a 30 MeV electron incident on gold, with emission at $\theta_{0}=0^{\circ}$. Bottom: Identical, for $\theta_{0}=5^{\circ}$.